# Realizable Continuous-Space Shields for Safe Reinforcement Learning

### v1

@@ Abstract
While Deep Reinforcement Learning (DRL) has achieved remarkable success across various domains, it remains vulnerable to occasional catastrophic failures without additional safeguards. One effective solution to prevent these failures is to use a shield that validates and adjusts the agent’s actions to ensure compliance with a provided set of safety specifications. For real-life robot domains, it is desirable to be able to define such safety specifications over continuous state and action spaces to accurately account for system dynamics and calculate new safe actions that minimally alter the agent’s output. In this paper, we propose the first shielding approach to automatically guarantee the realizability of safety requirements for continuous state and action spaces. Realizability is an essential property that confirms the shield will always be able to generate a safe action for any state in the environment. We formally prove that realizability can also be verified with a stateful shield, enabling the incorporation of non-Markovian safety requirements. Finally, we demonstrate the effectiveness of our approach in ensuring safety without compromising policy accuracy by applying it to a navigation problem and a multi-agent particle environment.
Validate
Action
Return Safe
Action
Proper Shield
Optimization
Criteria
Continuous
State Space
Real Time
Naïve Conditions
Solver (e.g., Z3)
Alshiekh et al. (2018)
This Work
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has achieved impressive results in a wide range of fields, demonstrating its potential to solve complex problems that were previously considered intractable. These include the mastery of large games like Go (Silver et al. 2016) and Dota 2 (Berner et al. 2019), as well as real-world applications ranging from healthcare (Pore et al. 2021) to autonomous driving (Tai, Paolo, and Liu 2017) and robotics (Aractingi et al. 2023). However, despite these noteworthy advancements, even the most sophisticated DRL algorithms (Schulman et al. 2017) face considerable challenges when analyzed on specific corner cases, where they persist in demonstrating a proclivity to commit critical mistakes (Corsi et al. 2024a; Szegedy et al. 2013). These limitations present a threat to the reliability of DRL systems, particularly when deployed in safety-critical applications, where even a single failure can have potentially catastrophic consequences (Srinivasan et al. 2020; Marvi and Kiumarsi 2021; Katz et al. 2019; Corsi, Marchesini, and Farinelli 2021).
@@
Traditional techniques to address safety concerns aim to embed this aspect as part of the learning process; some examples include reward shaping (Tessler, Mankowitz, and Mannor 2018), constrained reinforcement learning (Achiam et al. 2017; Ray, Achiam, and Amodei 2019), and adversarial training (Pinto et al. 2017). While these approaches can significantly improve the general reliability of the policy, they often provide only empirical guarantees, and the benefits are often presented in expectation (Srinivasan et al. 2020; He, Zhao, and Liu 2023). Although these may be sufficient for many problems, they cannot guarantee the safety of a system, which prevents their application in particularly safety-critical contexts.
@@
In contrast, a promising family of approaches provides formal safety guarantees through the adoption of an external component, commonly referred to as a shield (Garcıa and Fernández 2015; Corsi et al. 2024b). In DRL, a shield acts as a protective wrapper over the agent to ensure that its actions remain within safe boundaries, effectively preventing it from making dangerous or undesired decisions. However, most of the shielding techniques presented in the literature face a critical challenge: there are instances where no available action exists that satisfies all safety criteria simultaneously (Alshiekh et al. 2018). In practice, when the shielding mechanism encounters such scenarios, it fails to provide a safe action, leading to unsafe outcomes.
In this paper, we consider this crucial aspect of safety requirement satisfiability for all states. We define a Proper Shield as a shield that is always able to return an action that respects the given specifications for any state of the system (a formal definition is presented as Def. 1).
In the logic community, ensuring this property is equivalent to guaranteeing the realizability property for the shield (a formal definition is provided in section 3). However, it is particularly challenging for a human to provide a shield with safety requirements that it can always satisfy due to the compounding complexity of multiple safety requirements over continuous state and action spaces with many possible scenarios that can occur. Thus to prevent deploying a shield with unrealizable safety requirements, we propose an approach that includes an automatic realizability check as part of an offline procedure performed before deployment (an overview of the overall pipeline is provided in Fig.1). Intuitively, if the shield is not realizable, it means that the system cannot be safe in the required sense, and it is necessary to modify the dynamics of the system or relax the safety specifications. However, there is a scarcity of tools in the literature that support automatic realizability checking. A prominent example is the work of Alshiekh et al. (2018), which proposes to design a shield relying on the concept of Linear Temporal Logic (LTL). Their approach in particular allows the synthesis of the shield as a formal tool to automatically guarantee the realizability of the system. However, this technique comes with a crucial limitation: the synthesizer for the LTL specifications is designed to handle only discrete states and actions. This constraint is overly restrictive for real-world reinforcement learning applications such as robotics where the state and action spaces are often continuous and high dimensional.
This paper primarily builds upon the recent work on Realizability Modulo Theories of Rodríguez and Sánchez (2023). Applying their proposed satisfiability and realizability solver, we introduce a framework that for the first time is able to generate a Proper Shield capable of handling continuous state and action spaces. This significant advance allows us to augment a realizable shield with the benefits that come with using continuous state and action spaces. Firstly, our approach allows us to use optimization techniques to return actions that are not only safe but also closely aligned with the original decisions proposed by the neural network. This ensures that the agent’s behavior remains as optimal and effective as possible while adhering to safety constraints. Secondly, our method allows us to incorporate robot dynamics directly into the safety requirements, allowing for a more accurate model-based representation of the physical world in which the agent’s shield operates.
In addition, we formally prove that with our approach realizability can be verified with a stateful shield. This capability is crucial as it allows us to encode and enforce safety requirements that span multiple steps in the environment. Moreover, we demonstrate the effectiveness of our shielding approach in a mapless navigation domain, with additional experiments in a particle world multi-agent environment. We guarantee the realizability of safety requirements and ensure safety with continuous states and actions without compromising policy performance. To better understand the main contributions of our method and its position in existing literature, Table 1 presents a comparison with alternative techniques, highlighting the strengths of our approach compared to the baselines.
2 Related Work
Safe reinforcement learning aims to develop policies that not only optimize performance but also respect safety constraints, preventing the agent from taking actions that could lead to harmful or undesirable outcomes. In the literature, this safety aspect can be approached from many different perspectives. A first family of approaches attempts to enforce safer behavior as part of the learning process. An example is constrained reinforcement learning, where the MDP is enriched with a cost function to minimize that represents the safety requirements (He, Zhao, and Liu 2023; Achiam et al. 2017; Liu, Ding, and Liu 2020; Ray, Achiam, and Amodei 2019). Other algorithms focus on reward shaping approaches (Tessler, Mankowitz, and Mannor 2018), policy transfer (Simão, Tindemans, and Spaan 2022), behavioral monitors (Corsi et al. 2022; Srinivasan et al. 2020), or on techniques that can guarantee safe exploration and data collection phases (Marvi and Kiumarsi 2021). However, these approaches are not designed to provide formal guarantees about the final policy, especially in the context of deep learning where neural networks are known to be vulnerable to specific input configurations even when trained with state-of-the-art algorithms (Szegedy et al. 2013). Other approaches rely on verification methods to provide guarantees about the behavior of the neural network (Katz et al. 2019; Wang et al. 2021; Katz et al. 2017); however, they suffer from significant scalability problems and are limited in neural network topology (Liu et al. 2021). Moreover, as many recent papers have shown, as the number of requirements grows, it becomes nearly impossible to find models that formally respect the requirements over the entire input space (Corsi et al. 2024a; Amir et al. 2023).
@@ 3 Preliminaries
Reinforcement Learning (RL). We model interaction with an environment with a Partially Observable Markov Decision Process (POMDP), defined by a tuple $(X,A,O,P,R,\Omega,\gamma)$, with a state space of $X$, action space $A$, and observation space $O$. $P:X\times A\to\Delta(X)$ represents the state transition function. $R:X\times A\to\mathbb{R}$ is the reward function. $\Omega:X\to\Delta(O)$ represents the observation function, and $\gamma\in[0,1]$ is the discount factor. Our objective is to learn a policy $\pi:O\to\Delta(A)$ that maximizes the expected sum of discounted rewards over time. This is defined as $\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(x_{t},a_{t})\right]$, where $x_{t}$, $a_{t}$, and $o_{t}$ are the state, action, and observation at time $t$ until the episode horizon $T$.
Reactive Synthesis
@@
Satisfiability.
Given an LTL formula $\varphi$, the satisfiability problem determines if there is any possible execution that satisfies the specified temporal properties; i.e., we say $\varphi$ is satisfiable if there is a possible assignment of the variables in $\varphi$ such that $\varphi$ is satisfied. More formally, given variables $v=\{v^{0},v^{1},...\}$ in $\varphi$, if $\exists v$ s.t. $\varphi(v)$ holds, then $\varphi$ is satisfiable.
Realizability.
In reactive systems, the system must continuously interact with its environment and adapt its behavior accordingly. Thus, we need a property stronger than satisfiability: realizability, in which the variables of $\varphi$ are split into an uncontrollable player (i.e., the inputs provided by the environment) and a controllable player (i.e., the outputs provided by the system). Then, a formula $\varphi$ is realizable if for all possible valuations of the input variables, the output variables can be assigned so that the $\varphi$ is not violated. More formally, given environment variables $e^{0},e^{1},...$ and system variables $a^{0},a^{1},...$ of $\varphi$, if $(\forall e^{0},e^{1},...\exists a^{0},a^{1},...)^{*}$ s.t. $\varphi(e^{0},e^{1},...,a^{0},a^{1},...)$, then $\varphi$ is realizable [^1] Note that * is the Kleene closure (Lee 1982).. Moreover, if $\varphi$ is realizable, there is some strategy of the system: i.e., a way to assign the variables it controls such that $\varphi$ is never violated. The problem of computing such a strategy in realizable $\varphi$ is called reactive synthesis.
4 Continuous Shields
We now introduce our concept of a Proper Shield which is guaranteed to be realizable, providing safe alternatives to unsafe actions in all states. Working in a POMDP, we instantiate our shield as a stateful component where the shield’s state $h\in H$ is defined as the sequence of the constraints that held in the previous timesteps for an episode. To connect components in RL with the concepts in LTL, we consider the observation $o=\{o^{0},o^{1},...\},o\in O$ and $h\in H$ to be the uncontrollable variables and the action $a=\{a^{0},a^{1},...\},a\in A$ as the controllable variables.
Definition 1 (Proper Shield)
@@ 4 Continuous Shields
While we focus on the continuous case, this definition holds for both continuous and discrete spaces. In the following sections, we describe how we implement the offline realizability check and the online shielding process for continuous state and action spaces.
Architecture Overview
@@
In this paper we show how to use shield-synthesis, where reactive synthesis computes a shield $\zeta$ from $\varphi$ (in LTL modulo real arithmetic, i.e., continuous domains) and $\zeta$ is combined with an external agent $\pi$ (in our case, an RL-trained agent) to ensure that the composition $\pi\cdot\zeta$ will never violate $\varphi$. In each timestep, given the shield history $h$: (1) $\pi$ receives observation inputs $o$ from the environment and produces action outputs $a$; (2) $\zeta$ checks whether $\varphi(a,o,h)$ holds; and (3) if $\varphi(a,o,h)$ holds, the shield does not intervene, whereas in case of a violation, it overrides $a$ with a correction $\hat{a}=\zeta(a,o,h)$ such that $\varphi(\hat{a},o,h)$ now holds in that step.
This online process translates to providing a valid action (i.e., the assignment of the variables controlled by the system) that satisfies the specified requirements given the current state of the environment. The shield will propagate the continuous constraints through time and will check if the agent’s action satisfies all requirements and otherwise provide an alternative safe action. In LTLt, this is done by performing a query to a constraint satisfaction solver. If the constraint specification is not realizable, the shield may return an unsatisfiable (UNSAT) result, indicating that no safe action is possible given the current state.
In our framework (see Fig. 1(left)) we first check offline whether $\varphi$ is realizable, which ensures that for any possible state of the environment, the shield can always return a safe action. Moreover, since the continuous domain is a metric space, we can measure how far the correction is from the original candidate output. Therefore, we can also optimize the shield’s safe-action correction by returning the safe value closest to the agent’s proposed action.
Non-Markovian Requirements

### v2

@@ Abstract
While Deep Reinforcement Learning (DRL) has achieved remarkable success across various domains, it remains vulnerable to occasional catastrophic failures without additional safeguards. An effective solution to prevent these failures is to use a shield that validates and adjusts the agent’s actions to ensure compliance with a provided set of safety specifications. For real-world robotic domains, it is essential to define safety specifications over continuous state and action spaces to accurately account for system dynamics and compute new actions that minimally deviate from the agent’s original decision. In this paper, we present the first shielding approach specifically designed to ensure the satisfaction of safety requirements in continuous state and action spaces, making it suitable for practical robotic applications. Our method builds upon realizability, an essential property that confirms the shield will always be able to generate a safe action for any state in the environment. We formally prove that realizability can be verified for stateful shields, enabling the incorporation of non-Markovian safety requirements, such as loop avoidance. Finally, we demonstrate the effectiveness of our approach in ensuring safety without compromising the policy’s success rate by applying it to a navigation problem and a multi-agent particle environment
keywords:
Shielding, Reinforcement Learning, Safety, Robotics
@@ 1 Introduction
Deep Reinforcement Learning (DRL) has achieved impressive results in a wide range of fields, from mastering complex games like Go (Silver et al., 2016) and Dota 2 (Berner et al., 2019) to real-world applications in healthcare (Pore et al., 2021), autonomous driving (Tai et al., 2017), and robotics (Aractingi et al., 2023). However, even advanced DRL algorithms (Schulman et al., 2017) face considerable challenges when analyzed on specific corner cases, where they persist in demonstrating a proclivity to commit critical mistakes (Corsi et al., 2024a; Szegedy et al., 2013). Such limitations present a threat to the reliability of DRL systems, particularly when deployed in safety-critical applications, where even a single failure can have potentially catastrophic consequences (Srinivasan et al., 2020; Marvi and Kiumarsi, 2021; Katz et al., 2019; Corsi et al., 2021).
Traditional techniques to address safety concerns aim to embed this aspect as part of the learning process; some examples include sophisticated reward shaping techniques (Tessler et al., 2018), constrained reinforcement learning (Achiam et al., 2017; Ray et al., 2019), and adversarial training (Pinto et al., 2017). While these approaches can significantly enhance the overall reliability of the policy, their guarantees are typically empirical and their benefits are provided only in expectation (Srinivasan et al., 2020; He et al., 2023). Although these may be sufficient for many problems, they cannot guarantee the safety of a system, limiting their applicability in highly safety-critical contexts.
@@
To face this limitation, a promising family of approaches provides formal safety guarantees through the adoption of an external component, commonly referred to as a shield (Garcıa and Fernández, 2015; Corsi et al., 2024b). A shield acts as a protective wrapper over the agent to ensure that its actions remain within safe boundaries, effectively preventing it from making dangerous or undesired decisions (Alshiekh et al., 2018). However, most shielding techniques in the literature face a fundamental challenge: there are states where no action satisfies all safety criteria simultaneously. This issue is particularly critical in robotics. When an agent encounters a scenario where no safe action is available from the shield, it is left uncertain about how to proceed, as a default safe action is not always feasible in complex and dynamic environments. Therefore, it is essential to develop a shield that can always provide an action satisfying the given specifications for any state of the system, which we define as a Proper Shield.
In the logic community, ensuring this property is equivalent to guaranteeing the realizability of the shield. Alshiekh et al. (2018) proposed a method to design a proper shield using Linear Temporal Logic (LTL). Their approach uses the synthesis of the shield as a formal tool to automatically guarantee the realizability of the system. However, this method is limited to discrete states and actions, as the LTL synthesizer is not designed to handle continuous spaces. This constraint is overly restrictive for real-world reinforcement learning applications such as robotics where the state and action spaces are often continuous, high dimensional and nonlinear. More in general, despite its necessity and importance, little has been explored regarding continuous-space proper shields in robotics, Indeed, compared to discrete spaces, ensuring a proper shield in continuous spaces is significantly more difficult and remains an open challenge.
In this work, we extend a proper shield to continuous spaces, addressing a gap in existing methods. Our approach builds on the theoretical foundations provided by Rodríguez and Sánchez (2023), which solve the realizability problem for some fragments of LTLt, an extension of LTL that allows to include real values in the specification. While their work focuses purely on logical formulations with no direct connection to sequential decision-making, we adapt these ideas to develop a practical framework for generating a proper shield for continuous control tasks. As shown in Fig. 1, realizability checks are part of an offline procedure conducted prior to deployment. Crucially, working in continuous spaces allows us to encode a more accurate, model-based representation of the physical world within the specifications. Additionally, working with real values allows us to integrate optimization techniques to ensure that the shield not only returns safe actions but also provides decisions closely aligned with those proposed by the agent. This ensures the agent’s behavior remains both effective and optimal while adhering to strict safety constraints.
Additionally, we formally prove that realizability can be verified for shields with non-Markovian safety requirements. While the realizability of continuous non-Markovian requirements is typically undecidable, we address this challenge by introducing a so-called anticipation fragment of LTLt that eliminates operators requiring infinite memory, assuming deterministic dynamics. This fragment is expressive enough to capture relevant safety requirements under assumptions that are common in practical scenarios. For example, in our case study, we encode a rule to avoid loops within a specified time window (e.g., do not visit the same region for $i$ timesteps) to enhance success rates.
We demonstrate the effectiveness of our shielding approach in a mapless navigation domain, with additional experiments in a particle world multi-agent environment. In the mapless navigation domain, the shield effectively protects the agent from obstacles in an unknown environment, highlighting the shield’s ability to enforce safety without relying on pre-defined maps or global knowledge. We guarantee the realizability of safety requirements, ensuring that the agent has always at least one safe action available. Furthermore, with the integration of optimization techniques and the non-Markovian requirements we achieve the safety of the system with a minimal impact on the policy performance. In the particle-world multi-agent environment, our shield manages interactions among multiple agents in a shared, continuous space.
In summary, the contributions of this paper are as follows: (1) We introduce continuous-space proper shields for safe control. A detailed comparison with alternative shielding techniques is provided in Tab. 4 in the Appendix. (2) We propose a shield that enables formal realizability proofs for non-Markovian requirements by introducing the ‘anticipation fragment’ of LTLt, a decidable subset designed for stateful shields. This fragment effectively captures safety requirements under reasonable assumptions for practical use. (3) We demonstrate the merits of our method on robotic applications, including a mapless navigation domain and a particle-world multi-agent environment.
@@ 2 Preliminaries
We model interaction with an environment with a Partially Observable Markov Decision Process (POMDP), defined by a tuple $(X,A,O,P,R,\Omega,\gamma)$, with a state space of $X$, action space $A$, and observation space $O$. $P:X\times A\to\Delta(X)$ represents the state transition function. $R:X\times A\to\mathbb{R}$ is the reward function. $\Omega:X\to\Delta(O)$ represents the observation function, and $\gamma\in[0,1]$ is the discount factor. Our objective is to learn a policy $\pi:O\to\Delta(A)$ that maximizes the expected sum of discounted rewards over time. This is defined as $\mathbb{E}\left[\sum_{t=0}^{T}\gamma^{t}R(x_{t},a_{t})\right]$, where $x_{t}$, $a_{t}$, and $o_{t}$ are the state, action, and observation at time $t$ until the episode horizon $T$.
2.1 Concepts in Temporal Logic and Reactive Systems
@@
Satisfiability and Realizability.
Given an LTL formula $\varphi$, the satisfiability problem determines if there is any possible execution that satisfies the specified temporal properties; i.e., we say $\varphi$ is satisfiable if there is a possible assignment of the variables in $\varphi$ such that $\varphi$ is satisfied. More formally, given variables $v=\{v^{0},v^{1},...\}$ in $\varphi$, if $\exists v$ s.t. $\varphi(v)$ holds, then $\varphi$ is satisfiable. A reactive system is a type of system that continuously interacts with its environment by responding to inputs and adapting its behavior accordingly. To ensure such a system can meet the requirements under all possible conditions, we need a property stronger than satisfiability: realizability. In realizability, the variables of $\varphi$ are divided into an uncontrollable player (i.e., the inputs provided by the environment) and a controllable player (i.e., the outputs provided by the system). Then, a formula $\varphi$ is realizable if for all possible valuations of the input variables, the output variables can be assigned so that the $\varphi$ is not violated.
Realizability for Continuous Domains.
LTL is a powerful tool for reasoning about temporal properties. However, it lacks the expressiveness needed to handle continuous values, which are essential for realistic robotics applications to accurately encode environment dynamics in continuous spaces. To address this, we use Linear Temporal Logic modulo theories (LTLt), an extension of LTL that allows formulas to include variables from a first-order theory $\mathcal{T}$. While this has a precise meaning in formal verification  [^2] We invite the reader to read the supplementary material (Appendix E) for further insights., for simplicity in this paper, it means that LTLt enables specifying more complex properties involving both the temporal behavior of a system and real values data. A recent development by Rodríguez and Sánchez (2024b) solved the problem of realizability for certain fragments of LTLt. This makes it feasible to create realizable continuous-space shields using LTLt to ensure such expressive properties; for example, $(v^{1}>2.5)\rightarrow\ocircle(v^{2}>v^{1})$. Note that another powerful operator in LTLt is $\lhd v$, which, given $v$ allows access the value of $v$ in the previous timestep. However, this expressivity comes at the expense of the LTLt not being undecidable anymore (i.e., realizability checking procedures are not guaranteed to terminate).
3 Realizable Continuous Shields with Non-Markovian Requirements
A deep reinforcement learning shield is an external component that works on top of a policy to ensure compliance with a set of safety requirements $\varphi$. A Proper Shield is a special case of a shield that can provide safe alternatives to unsafe actions for all states in the system.
In this paper, to address the complexity of real-world robotic problems, our shield is designed to be stateful, enabling it to handle non-Markovian requirements. The shield maintains a state $h\in H$ in which, by construction, $h\in H$ represents only safe states (i.e., states where $\varphi$ has not been violated), as any violation in the past would contradict the safety specifications. Formally, given an action $a\in A$, an observation $o\in O$, a shield state $h\in H$, and a set of safety specifications $\varphi$, we define a Proper Shield as follows:
Definition 3.1 (Proper Shield).
@@ 3 Realizable Continuous Shields with Non-Markovian Requirements
Intuitively, a Proper Shield is a function $\zeta$ designed to always return an action $a$ such that $\varphi$ holds when starting from the initial state. The shield $\zeta$ is combined with an external agent $\pi$ (in our case, an RL agent) to ensure that the composition $\pi\cdot\zeta$ never violates $\varphi$. At each time step, given the shield’s history $h$: (1) $\pi$ receives observation inputs $o$ from the environment and produces an action $a$; (2) we check whether $\varphi(a,o,h)$ holds; and (3) if $\varphi(a,o,h)$ holds, the shield does not intervene, but if there is a violation, it overrides $a$ with a corrected action $\hat{a}=\zeta(a,o,h)$ such that $\varphi(\hat{a},o,h)$ now holds for that step, as described in Fig. 1 (right).
Ensuring that $\varphi$ is realizable is essential, as unrealizability could lead to situations where no safe action is possible for a given state. In such cases, computing $\hat{a}$ is not possible because the formula is unsatisfiable (UNSAT). In our framework (see Fig. 1, left), we check the realizability of $\varphi$ as part of an offline process prior to deployment, treating the observation $o$ and shield state $h$ as uncontrollable variables, while considering the action $a$ as a controllable variable.
Alshiekh et al. (2018) showed that a proper shield can be constructed using LTL, but their approach is limited to discrete state and action spaces. This restricts its applicability to robotics tasks and fails to capture accurate environment dynamics. In contrast, our shield leverages the more expressive specification language, LTLt, enabling the construction of a continuous-space proper shield that incorporates accurate dynamics directly into the specification. Moreover, since the continuous domain is a continuous space, we can minimize a distance metric between the correction $\hat{a}$ and the original candidate output $a$. This allows us to optimize the shield’s safe-action correction by returning $\hat{a}$ such that it is the closest safe value to the agent’s proposed action $a$.
3.1 Non-Markovian Encoding in LTLt
By exploiting recent advancements in LTLt synthesis as discussed in the previous sections, it becomes possible to encode Markovian properties that can be synthesized to create a proper shield:
Example 3.2.
Consider a simple agent that moves along a one-dimensional line, where its state is represented by a single variable $x$. The action space consists of a single action $a$, which specifies both the direction and magnitude of the step. The dynamics are given by $x_{t+1}=x_{t}+a_{t}$. Suppose we wish to enforce a simple safety requirement that ensures the agent never moves to a position lower than zero. This requirement can be encoded as follows:
$[x<1]\rightarrow\neg[a<-1]$
Note that this requirement is highly simplified; the challenge for the solver arises from the exponential growth in the number and complexity of such requirements. Nevertheless, properties of this form are often sufficient to encode essential safety requirements, such as collision avoidance. However, for real-world problems, we often need to incorporate requirements that extend beyond simple input-output relationships, involving multiple steps in the environment and sequences of the agent’s decisions. In such cases, it is necessary to encode non-Markovian requirements (or multistep properties) in a decidable fragment of LTLt, that guarantees an automatic realizability check. To better explain our approach we rely on a running example:
Example 3.3.
Let us revisit the agent from Example 3.2. We may now want to ensure that, once the agent visits a specific region of the state space, it will never return to the same position within a finite time horizon $i$. Specifically, we subdivide the state space in a finite number of intervals $r\in R$, where $r=[r_{l},r_{u}]$. This requirement requires a non-Markovian constraint as it involves consideration of multiple steps in the environment. We can formally encode the property using an LTLt formula $\varphi=A\rightarrow G$, where $A$ represents the assumptions, and $G$ specifies the guarantees the system must uphold under those assumptions: